# Smooth Tests for Normality in ANOVA

### v2

@@ 3.2 Different means and same variance
Next we allow the group means $\mu$ in model (3.1) to vary across the group index $j$, in the sense that
@@
where $\{\mu_{j}\}_{j=1}^{J}$ represents the potentially different means of each group. Let $\hat{\mu}_{j}=N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ and $\hat{\sigma}^{2}=N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\hat{\mu}_{j}\right)^{2}$ be estimators of $\mu_{j},j=1,\ldots,J$ and $\sigma^{2}$, respectively. Correspondingly we define $\hat{\varepsilon}_{ij}=Y_{ij}-\hat{\mu}_{j}$, $\hat{e}_{ij}=\hat{\varepsilon}_{ij}/\hat{\sigma}=(Y_{ij}-\hat{\mu}_{j})/\hat{\sigma}$, and $\hat{Z}_{ij}=\Phi(\hat{e}_{ij})$. The following theorem presents the null limiting properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ under the settings for model (3.9).
@@
as $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$ and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$. Furthermore, for $\hat{\Psi}_{K}^{2}$ defined in (3.3), we still have
@@
Theorem 4 is parallel with Theorem 1 and Corollary 1 except for extra conditions regarding the group sample sizes and the number of groups. These additional conditions are imposed for technical considerations and to account for the heterogeneity across group means from model (3.9). We illustrate the intuition behind the conditions as follows. First, we require that the sample size of each group, $N_{j}$, tends to infinity. For model (3.9) where $\mu_{j}$ is allowed to be distinct, only the observations in group $j$ can be used to estimate $\mu_{j}$, thus $N_{j}\to\infty$ is necessary for the consistency of $\hat{\mu}_{j}$. Second, the number of groups $J$ is required to be finite or to diverge no faster than $N^{1/2}$, and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$ must be satisfied jointly for $\{N_{j}\}_{j=1}^{J}$ and $J$, which rules out the case that the number of groups grows faster than the group sample sizes. In contrast, for model (3.1) in the previous section, only the total sample size $N\to\infty$ suffices to derive the asymptotic properties regardless of the magnitude of $N_{j}$s and $J$. This is because in model (3.1), all groups share the same mean $\mu$ and variance $\sigma^{2}$. Thus, the observations $\{Y_{ij}\}_{i=1,j=1}^{N_{j},J}$ are all i.i.d., and we can just pool them together for efficient estimation and inference.
The following two theorems establish the asymptotic properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ and $\hat{\Psi}_{K}^{2}$ (defined by (3.3)) under the alternatives. The conditions are the same as those of Theorem 4, and the results are parallel with those in Theorems 2 and 3.
@@ 3.3 Same mean and different variances
Now we allow the group variances $\sigma$ in model (3.1) to be different among the groups, in the sense that
@@
where $\{\sigma_{j}\}_{j=1}^{J}$ represents the potentially different standard deviations of each group. This time, we adopt the estimators $\hat{\mu}=J^{-1}\sum_{j=1}^{J}N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ and $\hat{\sigma}_{j}^{2}=N_{j}^{-1}\sum_{i=1}^{N_{j}}(Y_{ij}-\hat{\mu})^{2}$ for $\mu$ and $\sigma_{j}^{2},j=1,\ldots,J$, respectively. Note that the estimator of the population mean here is different from the sample mean in Section 3.1 due to the heterogeneity of variation across groups. Let $\hat{\varepsilon}_{ij}=Y_{ij}-\hat{\mu}$, $\hat{e}_{ij}=\hat{\varepsilon}_{ij}/\hat{\sigma}_{j}=(Y_{ij}-\hat{\mu})/\hat{\sigma}_{j}$ and $\hat{Z}_{ij}=\Phi(\hat{e}_{ij})$. Define $p_{j}=\lim N_{j}/N$ and $q_{j}=\lim JN_{j}/N$ regarding the relative proportion among the group sample sizes. Here we further impose the following conditions for model (3.10): (i) there exist $0<\underline{\sigma}\leq\overline{\sigma}<\infty$ such that $\underline{\sigma}<\inf_{1\leq j\leq J}\sigma_{j}\leq\sup_{1\leq j\leq J}\sigma_{j}<\overline{\sigma}$; (ii) there exist $0<\underline{q}\leq\overline{q}<\infty$ such that $\underline{q}<\inf_{1\leq j\leq J}q_{j}\leq\sup_{1\leq j\leq J}q_{j}<\overline{q}$. They both imply that the sample sizes and the standard deviations of each group are of the same magnitude.
Like the procedure in Sections 3.1 and 3.2 above, we start from the asymptotic behavior of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ here, see the following theorem.
@@
Note that the decomposition of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ in Theorem 7 is no longer the same as that in Theorems 1 and 4. The difference lies in the parameter estimation effect of $\hat{\mu}$ characterized by the term $c_{1k}N^{-1}(\sum_{\ell=1}^{J}p_{\ell}\sigma^{-1}_{\ell})\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\sigma_{j}e_{ij}/q_{j}$ on the right hand side of (3.11). Unlike the adoption of the sample mean for estimating $\mu$ in model (3.1), here we consider a new form of $\hat{\mu}$ for model (3.10), which leads to a different estimation effect term. Specially, when all groups share the same variance, i.e., $\sigma_{1}=\ldots=\sigma_{J}$, the term can be simplified as $c_{1k}N^{-1}\sum_{j=1}^{J}q_{j}^{-1}\sum_{i=1}^{N_{j}}e_{ij}=-c_{1k}\sum_{j=1}^{J}N_{j}^{-1}\sum_{i=1}^{N_{j}}e_{ij}$; further if the group sample sizes are equal, i.e., $N_{1}=\ldots=N_{J}$, then the estimation effect reduces to that in Theorem 1, $c_{1k}N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$. In contrast, the estimation effect of the variances is still aggregated as the term $c_{2k}(2N)^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)$ regardless of the equality of group variances or group sample sizes.
The asymptotic normality of the vector $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}(\hat{Z}_{ij})$ can be derived from (3.11). Under the conditions of Theorem 7, by CLT,
@@
By replacing $\sigma_{j}$, $p_{j}$ and $q_{j}$ by $\hat{\sigma}_{j}$, $\hat{p}_{j}=N_{j}/N$ and $\hat{q}_{j}=JN_{j}/N$ respectively, we obtain a consistent estimator of $\bm{\Omega}_{K}^{(j)}$, denoted as $\widehat{\bm{\Omega}}_{K}^{(j)}$. Correspondingly, the test statistic takes the form
@@
We also define the infeasible version,
@@
Corollary 2 demonstrates that under the null hypothesis, both the feasible test statistic $\hat{\Psi}_{K}^{2}$ (in (3.12)) and its infeasible version $\tilde{\Psi}_{K}^{2}$ (in (3.13)) are asymptotically $\chi_{K}^{2}$ distributed after multiplied by $N$, the total sample size. Although the test statistics considered here take a slightly different form from those in the earlier setting, the limiting Chi-square distribution is retained and the convergence rate remains $N$, thereby reinforcing the validity of our unified theoretical framework. The following two theorems establish the asymptotic behavior of $\tilde{\Psi}_{K}^{2}$ under the fixed and local alternatives, which is also analogous to that of $\hat{\Psi}_{K}^{2}$ in Sections 3.1 and 3.2.
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{(\pi_{k}(Z_{ij})-\mathrm{E}\left[\pi_{k}(Z)\right])-d_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e_{ij}}{q_{j}}-\frac{d_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$
$+\mathrm{E}\left[\pi_{k}(Z)\right]+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (3.14)
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{(\pi_{k}(Z_{ij})-\mathrm{E}\left[\pi_{k}(Z)\right])-c_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e_{ij}}{q_{j}}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$
$+\delta_{N}\Delta_{k}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
@@
To conclude this section, we summarize the regularity conditions of the three models above in Table 1. On the one hand, one can find that more complex model structures lead to more restrictions on parameters and sample sizes. The heterogeneity across groups indeed poses additional challenges for theoretical analysis. On the other hand, under these mild conditions, an asymptotic Chi-square test with convergence rate $N$ can always be established for each of the three cases. This demonstrates that our proposed method possesses a unified theoretical framework while retaining the ability to capture the distinctive features of different ANOVA models.
@@ 4 Data-driven choice of $K$
The methodology in Section 3 is based on the fact that the order of smooth alternatives (2.1), $K$, is a predetermined fixed integer before the tests are performed. In fact, the order $K$ has a key impact on the performance of smooth tests (Ledwina,, 1994; Kallenberg and Ledwina, 1995a, ; Kallenberg and Ledwina, 1995b, ; Inglot and Ledwina,, 1996; Inglot et al.,, 1997; Kallenberg and Ledwina,, 1997; Kallenberg and Teresa,, 1997; Kallenberg and Ledwina,, 1999; Janic-Wróblewska and Ledwina,, 2000; Ducharme and Fontez,, 2004; Ducharme and Lafaye de Micheaux,, 2004; Kraus,, 2007; Duchesne et al.,, 2016). If $K$ is large, the alternatives we test against contain redundant components that do not contribute to the test statistic markedly but increase the number of degrees of freedom. This will make smooth tests more omnibus and lead to power dilution. If $K$ is small, as described in Section 3, it is more likely to happen that $\mathrm{E}[\pi_{k}(Z)]=0$ for all $1\leq k\leq K$, which makes the tests exhibit no power. Therefore, it is crucial to select an appropriate $K$ before applying the test. In this section, we focus on the data-driven choice of $K$ for the three cases in Section 3.
In the literature of Neyman’s smooth tests, the spirit of data-driven methods was first introduced by Ledwina, (1994) for the problem of testing for uniformity. The data-driven procedure consists of two steps. First, Schwarz’s selection rule (also known as the Bayesian Information Criterion, BIC) is applied to find a suitable dimension of the smooth model for the data. Second, Neyman’s smooth test is performed within the selected model space, yielding a data-driven test statistic. Therefore, the data-driven smooth test serves as a kind of first selection followed by a more precise instrument, being Neyman’s test in the “right” direction. Desirable theoretical properties and extensive numerical results of the selection rule and the induced test statistic were shown in Kallenberg and Ledwina, 1995a and Kallenberg and Ledwina, 1995b , Inglot and Ledwina, (1996). Later, some modified versions of Schwarz’s selection rule were employed to address more complicated issues, such as testing for composite hypotheses (Kallenberg and Ledwina,, 1997; Kallenberg and Teresa,, 1997; Inglot et al.,, 1997), testing for independence (Kallenberg and Ledwina,, 1999), two-sample testing problems (Janic-Wróblewska and Ledwina,, 2000), and so on. The modified selection rules utilize the smooth test statistics and circumvent the need to compute the maximized log-likelihood function of smooth models; thus, they are easier to implement.
Motivated by the strategies mentioned above, we adopt a modified Schwarz’s rule to determine the order $K$ for our problem. The selection rule is specified as
@@
where $D$ is some fixed positive integer, and $\hat{\Psi}_{k}^{2}$ takes the form of (3.3) or (3.12) up to specific scenarios. The corresponding data-driven test statistic is $N\Psi_{\hat{K}}^{2}$, where $\hat{K}$ is given by (4.1). The form of (4.1) is mainly inspired by Ducharme and Fontez, (2004), Ducharme and Lafaye de Micheaux, (2004), Kraus, (2007) and Duchesne et al., (2016) where the upper bound of selection $D$ is fixed. An alternative is to set $D=D(N)\to\infty$ as $N\to\infty$ (see, for example, Kallenberg and Ledwina, 1995a, ; Kallenberg and Ledwina, 1995b, ; Inglot and Ledwina,, 1996; Inglot et al.,, 1997; Kallenberg and Ledwina,, 1997; Kallenberg and Teresa,, 1997; Kallenberg and Ledwina,, 1999; Janic-Wróblewska and Ledwina,, 2000). Although the diverging upper bound helps to improve the consistency of smooth tests theoretically, the rate of divergence is required to be quite slow, and the empirical power levels off rapidly as $D$ increases in their simulation studies. Therefore, our proposed selection rule (4.1) with fixed $D$ is reasonable and simple to implement in practice.
To establish the properties of $\hat{K}$ and $N\Psi_{\hat{K}}^{2}$, we need a revised version of the alternative hypotheses $H_{1}$ (3.4) and Assumption 3.2. Consider the following alternatives:
$H_{1}^{\prime}:\mathrm{E}[\pi_{1}(Z)]=\ldots=\mathrm{E}[\pi_{K_{0}-1}(Z)]=0,\mathrm{E}[\pi_{K_{0}}(Z)]\neq 0,\quad K_{0}\leq D.$ (4.2)
In other words, $H_{1}^{\prime}$ encompasses a broader range of alternatives than $H_{1}$ due to the replacement of the fixed order $K$ with the larger $D$. Analogously, the condition on the orthonormal system is now broadened to include all functions up to order $D$; see the revised assumption below.
@@
The following theorem demonstrates the properties of data-driven tests under the null and the revised alternatives.
@@
Theorem 10 presents the unified results of data-driven smooth tests for the three cases in Section 3. On the one hand, under the null hypothesis, the probability of $\{\hat{K}=1\}$ tends to $1$ asymptotically, which implies that the first-order smooth model fits $\{\hat{Z}_{ij}\}_{i=1,j=1}^{N_{j}.J}$ best among the $D$ candidate models. From the perspective of hypothesis testing, it is $\hat{\Psi}_{1}^{2}$ that is informative and adequate for diagnosing the uniformity/normality. On the other hand, the data-driven selection procedure helps identify the mechanism behind nonuniformity/nonnormality and leads to a more powerful test against a broader range of alternatives, $H_{1}^{\prime}$. These findings are parallel with existing results about data-driven smooth tests.
However, in practice, the $\chi_{1}^{2}$ limiting null distribution of data-driven smooth test statistics is often inaccurate in finite samples (Kallenberg and Ledwina, 1995b, ; Inglot et al.,, 1997; Kallenberg and Ledwina,, 1997; Kallenberg and Teresa,, 1997; Kallenberg and Ledwina,, 1999; Janic-Wróblewska and Ledwina,, 2000; Ducharme and Fontez,, 2004; Ducharme and Lafaye de Micheaux,, 2004; Kraus,, 2007; Duchesne et al.,, 2016). To overcome this issue in our problem, provided the nested orthonormal system (i.e. $\{\pi_{k}\}_{k=1}^{K}\subseteq\{\pi_{k}\}_{k=1}^{K+1}$ for $1\leq K\leq D-1$), we follow the instruction of Kallenberg and Ledwina, 1995b , Inglot et al., (1997), Kallenberg and Ledwina, (1997), Kallenberg and Ledwina, (1999), Janic-Wróblewska and Ledwina, (2000), Kraus, (2007) and adopt a finite-sample approximation for the null distribution $N\hat{\Psi}_{\hat{K}}^{2}$:

### v3

@@ 3.2 Different means and same variance
Next, we consider ANOVA models in which the group means are allowed to differ, that is,
@@
where $\{\mu_{j}\}_{j=1}^{J}$ denote the potentially distinct means of each group. Let $\hat{\mu}_{j}=N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ and $\hat{\sigma}^{2}=N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\hat{\mu}_{j}\right)^{2}$ be the estimators of $\mu_{j}$ ($j=1,\ldots,J$) and $\sigma^{2}$, respectively. Correspondingly we define $\hat{\varepsilon}_{ij}=Y_{ij}-\hat{\mu}_{j}$, $\hat{e}_{ij}=\hat{\varepsilon}_{ij}/\hat{\sigma}=(Y_{ij}-\hat{\mu}_{j})/\hat{\sigma}$, and $\hat{Z}_{ij}=\Phi(\hat{e}_{ij})$. The following theorem characterizes the properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ under the model (3.9).
@@
as $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$ and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$. Furthermore, for $\hat{\Psi}_{K}^{2}$ defined in (3.3), we have
@@
Theorem 4 parallels Theorem 1 and Corollary 1, except that it requires additional conditions on the group sample sizes and the number of groups. These conditions are imposed for technical reasons and to account for heterogeneity across group means in model (3.9). The intuition behind the conditions is illustrated as follows. First, the sample size of each group, $N_{j}$, must tend to infinity. In model (3.9), where $\mu_{j}$ can be distinct, only the observations within group $j$ can be used to estimate $\mu_{j}$, making $N_{j}\to\infty$ necessary for the consistency of $\hat{\mu}_{j}$. Second, the number of groups $J$ is required to be finite or to diverge no faster than $N^{1/2}$, and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$ must hold jointly for $\{N_{j}\}_{j=1}^{J}$ and $J$. This rules out scenarios where the number of groups grows faster than the sample sizes within each group. In contrast, for model (3.1) in the previous section, it suffices that the total sample size $N\to\infty$ to derive asymptotic properties, regardless of the magnitudes of $N_{j}$ and $J$. This is because in model (3.1), all groups share the same mean $\mu$ and variance $\sigma^{2}$, so the observations $\{Y_{ij}\}_{i=1,j=1}^{N_{j},J}$ are i.i.d., allowing them to be pooled for efficient estimation and inference.
The following two theorems establish formally the asymptotic properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ and $\hat{\Psi}_{K}^{2}$ (defined in (3.3)) under the alternatives. The conditions are identical to those in Theorem 4, and the results parallel those in Theorems 2 and 3.
@@ 3.3 Same mean and different variances
We now extend model (3.1) by allowing the group variances to differ across groups, that is,
@@
where $\{\sigma_{j}\}_{j=1}^{J}$ represents the potentially distinct standard deviations of each group. For estimation, we adopt $\hat{\mu}=J^{-1}\sum_{j=1}^{J}N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ and $\hat{\sigma}_{j}^{2}=N_{j}^{-1}\sum_{i=1}^{N_{j}}(Y_{ij}-\hat{\mu})^{2}$ for $\mu$ and $\sigma_{j}^{2}$ ($j=1,\ldots,J$), respectively. Note that the estimator of the population mean here is different from the sample mean in Section 3.1 due to the heterogeneity of group variances. Define $\hat{\varepsilon}_{ij}=Y_{ij}-\hat{\mu}$, $\hat{e}_{ij}=\hat{\varepsilon}_{ij}/\hat{\sigma}_{j}=(Y_{ij}-\hat{\mu})/\hat{\sigma}_{j}$ and $\hat{Z}_{ij}=\Phi(\hat{e}_{ij})$. Let $p_{j}=\lim N_{j}/N$ and $q_{j}=\lim JN_{j}/N$ be quantities characterizing the relative proportions of the group sample sizes. Here we further impose the following conditions for model (3.10): (i) there exist $0<\underline{\sigma}\leq\overline{\sigma}<\infty$ such that $\underline{\sigma}<\inf_{1\leq j\leq J}\sigma_{j}\leq\sup_{1\leq j\leq J}\sigma_{j}<\overline{\sigma}$; (ii) there exist $0<\underline{q}\leq\overline{q}<\infty$ such that $\underline{q}<\inf_{1\leq j\leq J}q_{j}\leq\sup_{1\leq j\leq J}q_{j}<\overline{q}$. These conditions imply that the group-specific sample sizes and standard deviations are of the same order of magnitude.
Following the approaches in Sections 3.1 and 3.2, we first examine the asymptotic behavior of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$, which is summarized in the theorem below.
@@
Note that the asymptotic decomposition of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ in Theorem 7 is no longer the same as that in Theorems 1 and 4. The difference lies in the parameter estimation effect of $\hat{\mu}$, which is reflected by the term $c_{1k}N^{-1}(\sum_{\ell=1}^{J}p_{\ell}\sigma^{-1}_{\ell})\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\sigma_{j}e_{ij}/q_{j}$ on the right-hand side of (3.11). Unlike model (3.1), which uses the sample mean to estimate $\mu$, model (3.10) employs a different estimator, $\hat{\mu}$, resulting in this distinct estimation effect. In particular, when all groups share the same variance, i.e., $\sigma_{1}=\ldots=\sigma_{J}$, the term reduces to $c_{1k}N^{-1}\sum_{j=1}^{J}q_{j}^{-1}\sum_{i=1}^{N_{j}}e_{ij}=-c_{1k}\sum_{j=1}^{J}N_{j}^{-1}\sum_{i=1}^{N_{j}}e_{ij}$; further if the group sample sizes are equal, i.e., $N_{1}=\ldots=N_{J}$, then the estimation effect simplifies to that in Theorem 1, $c_{1k}N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$. In contrast, the estimation effect of the variances always aggregates into the term $c_{2k}(2N)^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)$ regardless of the equality of group variances or group sample sizes.
The asymptotic normality of the vector $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}(\hat{Z}_{ij})$ follows directly from (3.11). Under the conditions of Theorem 7, an application of the CLT yields that
@@
By replacing $\sigma_{j}$, $p_{j}$ and $q_{j}$ with their sample analogues $\hat{\sigma}_{j}$, $\hat{p}_{j}=N_{j}/N$ and $\hat{q}_{j}=JN_{j}/N$ respectively, we obtain a consistent estimator of $\bm{\Omega}_{K}^{(j)}$, denoted by $\widehat{\bm{\Omega}}_{K}^{(j)}$. Accordingly, the test statistic is defined as
@@
We also introduce the infeasible version,
@@
Corollary 2 shows that under the null hypothesis, both the feasible test statistic $\hat{\Psi}_{K}^{2}$ (defined in (3.12)) and its infeasible counterpart $\tilde{\Psi}_{K}^{2}$ (in (3.13)) are asymptotically $\chi_{K}^{2}$ distributed when multiplied by the total sample size $N$. Although these test statistics take a slightly different form from those in the previous Sections 3.1 and 3.2, the limiting Chi-square distribution is retained, and the convergence rate remains $N$, which further supports the validity of our unified theoretical framework. The following two theorems characterize the asymptotic behavior of $\tilde{\Psi}_{K}^{2}$ under both fixed and local alternatives, in a manner analogous to that of $\hat{\Psi}_{K}^{2}$ discussed in Sections 3.1 and 3.2.
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathrm{E}\left[\pi_{k}(Z)\right]\right)-d_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e_{ij}}{q_{j}}\right$.
$-\left.\frac{d_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}+\mathrm{E}\left[\pi_{k}(Z)\right]+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (3.14)
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathrm{E}\left[\pi_{k}(Z)\right]\right)-c_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e_{ij}}{q_{j}}\right$.
$-\left.\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}+\delta_{N}\Delta_{k}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
@@
To conclude this section, we summarize the regularity conditions for the three models in Table 1. On one hand, more complex model structures naturally impose additional restrictions on parameters and group sample sizes, as the heterogeneity across groups introduces extra challenges for theoretical analysis. On the other hand, under these mild conditions, an asymptotic Chi-square test with convergence rate $N$ can be established for each model. This highlights that our proposed methodology provides a unified theoretical framework while remaining flexible enough to accommodate the distinctive features of different ANOVA settings.
@@ 4 Data-driven choice of $K$
The methodology in Section 3 relies on the fact that the order of smooth alternatives (2.1), $K$, is fixed before testing. The choice of $K$ critically affects the performance of smooth tests (Ledwina,, 1994; Kallenberg and Ledwina, 1995a, ; Kallenberg and Ledwina, 1995b, ; Inglot and Ledwina,, 1996; Inglot et al.,, 1997; Kallenberg and Ledwina,, 1997; Kallenberg and Teresa,, 1997; Kallenberg and Ledwina,, 1999; Janic-Wróblewska and Ledwina,, 2000; Ducharme and Fontez,, 2004; Ducharme and Lafaye de Micheaux,, 2004; Kraus,, 2007; Duchesne et al.,, 2016). A large $K$ introduces redundant components in the alternatives that contribute little to the test statistic but inflate the degrees of freedom, leading to power dilution. Conversely, a small $K$ increases the risk that $\mathrm{E}[\pi_{k}(Z)]=0$ for all $1\leq k\leq K$, rendering the test powerless. Hence, an appropriate choice of $K$ is crucial. In this section, we focus on data-driven selection of $K$ for the three cases in Section 3.
In the literature of Neyman’s smooth tests, data-driven selection was first proposed by Ledwina, (1994) for testing uniformity. The data-driven procedure consists of two steps. First, Schwarz’s selection rule (also known as the Bayesian Information Criterion, BIC) is applied to determine a suitable dimension for the smooth model that best fits the data. Second, Neyman’s smooth test is performed within the selected model space, yielding a data-driven test statistic. The data-driven smooth test thus serves to combine preliminary model selection with a more precise inferential procedure, being Neyman’s test in the “right” direction. Desirable theoretical properties and extensive numerical results of the selection rule and the induced test statistic were established in Kallenberg and Ledwina, 1995a and Kallenberg and Ledwina, 1995b , Inglot and Ledwina, (1996). Later, modifications of Schwarz’s rule addressed more complex problems, including testing for composite hypotheses (Kallenberg and Ledwina,, 1997; Kallenberg and Teresa,, 1997; Inglot et al.,, 1997), testing for independence (Kallenberg and Ledwina,, 1999), two-sample testing problems (Janic-Wróblewska and Ledwina,, 2000), and so on. These modified rules leverage smooth test statistics directly, avoiding the computation of the maximized log-likelihood, and are easier to implement.
Motivated by these strategies, we adopt a modified Schwarz’s rule to determine the order $K$ for smooth tests in ANOVA models. The selection rule is specified as follows:
@@
where $D$ is a fixed positive integer, and $\hat{\Psi}_{k}^{2}$ takes the form of (3.3) or (3.12) depending on the specific scenario. The resulting data-driven test statistic is $N\Psi_{\hat{K}}^{2}$, with $\hat{K}$ given by (4.1). The form of (4.1) is mainly inspired by Ducharme and Fontez, (2004), Ducharme and Lafaye de Micheaux, (2004), Kraus, (2007) and Duchesne et al., (2016) where the upper bound of selection $D$ is fixed. An alternative is to let $D=D(N)\to\infty$ as $N\to\infty$ (see, for example, Kallenberg and Ledwina, 1995a, ; Kallenberg and Ledwina, 1995b, ; Inglot and Ledwina,, 1996; Inglot et al.,, 1997; Kallenberg and Ledwina,, 1997; Kallenberg and Teresa,, 1997; Kallenberg and Ledwina,, 1999; Janic-Wróblewska and Ledwina,, 2000). Although a diverging upper bound can theoretically improve the consistency of smooth tests, the required rate of divergence is slow, and simulation studies show that empirical power levels off rapidly as $D$ increases. Hence, our proposed selection rule (4.1) with fixed $D$ is both reasonable and easy to implement in practice.
To establish the theoretical properties of $\hat{K}$ and $N\hat{\Psi}_{\hat{K}}^{2}$, we introduce a revised version of the alternative hypothesis $H_{1}$ (3.4) and Assumption 3.2. Consider
$H_{1}^{\prime}:\mathrm{E}[\pi_{1}(Z)]=\ldots=\mathrm{E}[\pi_{K_{0}-1}(Z)]=0,\mathrm{E}[\pi_{K_{0}}(Z)]\neq 0,\quad K_{0}\leq D,$ (4.2)
which encompasses a broader range of alternatives than $H_{1}$ by replacing the fixed order $K$ with the larger $D$. Correspondingly, the condition on the orthonormal system is extended to include all functions up to order $D$, as follows.
@@
The following theorem establishes the properties of data-driven tests under the null and the revised alternatives.
@@
Theorem 10 presents the unified results of data-driven smooth tests for the three cases in Section 3. Under the null hypothesis, the probability of $\{\hat{K}=1\}$ tends to $1$ asymptotically, implying that the first-order smooth model provides the best fit to $\{\hat{Z}_{ij}\}_{i=1,j=1}^{N_{j},J}$ among the $D$ candidate models. From the perspective of hypothesis testing, this means that $\hat{\Psi}_{1}^{2}$ is informative and sufficient for assessing uniformity or normality. Meanwhile, the data-driven selection procedure allows for the identification of the mechanism behind nonuniformity (and also nonnormality) and enhances the test’s power against a broader class of alternatives, $H_{1}^{\prime}$. These findings are consistent with previous results on data-driven smooth tests.
In practice, however, the $\chi_{1}^{2}$ limiting null distribution of data-driven smooth test statistics often performs poorly in finite samples (Kallenberg and Ledwina, 1995b, ; Inglot et al.,, 1997; Kallenberg and Ledwina,, 1997; Kallenberg and Teresa,, 1997; Kallenberg and Ledwina,, 1999; Janic-Wróblewska and Ledwina,, 2000; Ducharme and Fontez,, 2004; Ducharme and Lafaye de Micheaux,, 2004; Kraus,, 2007; Duchesne et al.,, 2016). To address this issue in our problem, and provided the nested orthonormal system (i.e. $\{\pi_{k}\}_{k=1}^{K}\subseteq\{\pi_{k}\}_{k=1}^{K+1}$ for $1\leq K\leq D-1$), we follow the approach of Kallenberg and Ledwina, 1995b , Inglot et al., (1997), Kallenberg and Ledwina, (1997), Kallenberg and Ledwina, (1999), Janic-Wróblewska and Ledwina, (2000), Kraus, (2007) and adopt the following finite-sample approximation for the null distribution of $N\hat{\Psi}_{\hat{K}}^{2}$: